Solve each system by substitution or addition, whichever is easier.
52. 2y - x = 3
x = 3y -5

Answer by ewatrrr(24785)


Hi
2y - x = 3
x = 3y -5 |using substitution to solve
2y - (3y-5) = 3
-y = -2
y = 2 and x = 1 ( x = 3*2-5)
CHECKING our Answer***
2y - x = 3
4 - 1 = 3